need to know what x equals
3/7 x + 5/14 = 19/42

Answer by rfer(16322)

3/7x=19/42-5/14
3/7x=19/42-15/42
3/7x=4/12
x=(4/12)/(3/7)
x=(4/12)(7/3)
x=28/36
x=7/9